OpenStudy (anonymous):

isoltating the variable 2b-9=1

OpenStudy (anonymous):

2b - 9 = 1 2b = 1 + 9 2b = 10 divide by 2: b = 5

OpenStudy (anonymous):

x-3=11

OpenStudy (anonymous):

Try it yourself, start with adding 3 to both sides.

OpenStudy (anonymous):

14?

OpenStudy (anonymous):

and 6

OpenStudy (anonymous):

Yeah x = 14.

OpenStudy (anonymous):

-5 + 1/5 a=4

OpenStudy (anonymous):

Do you mean \[-5 + \frac{ 1 }{ 5 }a = 4\]

OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

start with adding something to the equation to leave the variable alone.

OpenStudy (anonymous):

like what?

OpenStudy (anonymous):

what number will make 1/5 a stay alone on the left side?

OpenStudy (anonymous):

i took away 4 from -5 and 4

OpenStudy (anonymous):

So now you've got \[-9 + \frac{ 1 }{ 5 }a = 0\]

OpenStudy (anonymous):

oh like stay alone wait so i take away -5?

OpenStudy (anonymous):

That's right.

OpenStudy (anonymous):

ok so then -5-4?

OpenStudy (skullpatrol):

Do you know what an "inverse" operation is? @meggy

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

No, if you subtruct -5 it's actually like adding 5. -(-5) = 5

OpenStudy (skullpatrol):

Inverse operations: Operations that "undo" each other. For example multiplication and division.

OpenStudy (skullpatrol):

Do you understand that definition @meggy ?